# Surface hopping molecular dynamics simulation of ultrafast methyl iodide photodissociation mapped by Coulomb explosion imaging

### v1

@@ 3.4 Coulomb explosion at short pump-probe delays
Fig.  6: Simulated kinetic energy release at different pump-probe delays ranging from 0 fs to 50 fs using about 20,000 PD-CE trajectories at T=100 K. The Coulomb explosion is simulated on the lowest ionic potential energy surface of $\text{CH}_{3}\text{I}^{2+}$ that corresponds to $\text{I}^{+}(^{3}\text{P}_{2})$ final dissociation limit. (a) Signals including all active states during photodissociation, which will lead to both I and $\text{I}^{*}$ channels. (b) Signals only from 7th adiabatic state during photodissocaition, which correspond to I channel. (c) Signals only from 9th adiabatic state during photodissociation, which corresponds to $\text{I}^{*}$ channel at long distance. All signals are counted using a bin size $h_{t}=0.5$ fs and $h_{E}=0.01$ eV.
@@
At short pump-probe delays (t<50 fs), the $\text{CH}_{3}\text{I}$ molecule is in the intermediate states of photodissociation. Non-adiabatic transitions also take place at this time scale. To investigate the non-adiabatic dynamics in this regime, additional 20,000 PD-CE trajectories are calculated with initial conditions sampled at T=100 K and with a pump-probe delay step $\Delta t=0.1$ fs up to 50 fs. Here we choose a lower effective temperature to make the initial conditions a little more condensed in the phase space such that the non-adiabatic features are better resolved in KER signals. The Coulomb explosion part is still simulated on the lowest PES corresponding to $\text{I}^{+}(^{3}\text{P}_{2})$ threshold. The resulting delay-dependent KER signals are shown in Fig. 6. A clear bifurcation of signals is present between 10 fs and 20 fs in Fig. 6(a), which indicates the occurrence of non-adiabatic transitions. The photodissociation separates to two channels leading to I and I* products. Fig. 6(b) shows that only a small fraction of trajectories (about 20 percent) undergoes adiabatic passage and dissociate to I channel. Most trajectories (about 80 percent) start to hop to the 9th adiabatic state near 12 fs and dissociate to $\text{I}^{*}$ channel, as shown in Fig. 6(c). It should be noted that the 1D model is a poor approximation at short pump-probe delays. The C-I stretching is highly coupled with $\text{CH}_{3}$($\text{CH}_{3}^{+}$) umbrella motion and the C-I bending at small C-I distances, which is not characterized by the 1D model. Not only does the 1D model predict a fake potential well on $\text{I}^{+}(^{3}\text{P}_{2})$ ionic PES, but also it significantly underestimates the KER. Therefore, we do not compare our simulation with the 1D model at short pump-probe delays.
@@
The bifurcation of KER signals has not been observed in CEI experiments so far. One reason is that the conical intersection is very close to the FC point and $\text{CH}_{3}$ is a very light fragment. Thus, non-adiabatic transitions occur within 20 fs after UV excitation, which requires pump and probe pulses with very high temporal resolutions to resolve. One option to alleviate this difficulty might be choosing a molecule with heavier fragments such as $\text{CF}_{3}\text{I}$. An alternative method to map the non-adiabatic transitions is using ATAS, which has been studied both theoretically and experimentally55, 32.
@@
Fig.  7: Simulated $\text{CH}_{3}\text{I}^{2+}$ signals at very short pump-probe delays. (a) Potential energy surface of $\text{CH}_{3}\text{I}^{2+}$ corresponding to $\text{I}^{+}(^{1}\text{S}_{0})$ threshold, projected to $\alpha$ and $R$ plane. The energy difference between adjacent contours is 0.1 eV. The Coulomb explosion is simulated on this surface. 4,000 PD-CE trajectories are calculated at $T=100$ K with pump-probe delays from 0 fs to 10 fs. The black triangles, which are counted as $\text{CH}_{3}\text{I}^{2+}$ signals, indicate the final geometries of 956 PD-CE trajectories after propagating for over 2 ps on the ionic potential energy surface. (b) $\text{CH}_{3}\text{I}^{2+}$ signals as a function of pump-probe delay.
@@
As mentioned before, only the ionic state associated with $\text{I}^{+}(^{1}\text{S}_{0})$ threshold has a shallow well on the PES and can support a metastable $\text{CH}_{3}\text{I}^{2+}$ state, as shown in Fig. 7(a). But it should be noted that this channel has a rather small transition probability during the IR multi-photon ionization. In order to investigate the $\text{CH}_{3}\text{I}^{2+}$ signals, we calculate 4,000 PD-CE trajectories at very short pump-probe delays (t<10 fs), for which the Coulomb explosion is simulated on the $\text{I}^{+}(^{1}\text{S}_{0})$ ionic PES. 956 trajectories form $\text{CH}_{3}\text{I}^{2+}$ states and never break up into ionic fragments, as shown on top of the PES in Fig. 7(a). The $\text{CH}_{3}\text{I}^{2+}$ signal has a maximum yield near 1 fs, and damps rapidly with the pump-probe delay, as shown in Fig. 7(b). Due to this shallow well on the ionic PES, the $\text{CH}_{3}\text{I}^{2+}$ signals are detected in the CEI experiment27, and an enhancement of $\text{CH}_{3}\text{I}^{2+}$ yield is observed near zero pump-probe delay, which is attributed to the cross correlation between the pump and the probe pulses.
@@ 4 Conclusion and outlook
In this work, we have developed an effective method to directly simulate the photodissociation followed by Coulomb explosion of methyl iodide with the inclusion of non-adiabatic effects. Potential energy surfaces of both the neutral molecule and the doubly charged cation are built upon high-level ab initio electronic structure calculations. More than 40,000 trajectories are calculated on both the neutral and ionic PESs to obtain a statistically reliable interpretation of the delay-dependent KER signals. The non-adiabatic effects are treated using a Landau-Zener surface hopping algorithm, which yields a branching ratio in overall agreement with previous theories and experiments. Our simulation of the delay-dependent KER at large pump-probe delays shows good agreement with a recent experiment using coincident ion momentum imaging techniques27, with the KER signals reflecting two reaction channels. At short pump-probe delays, the simulation shows a clear bifurcation of the KER signals near 12 fs, indicating the occurrence of non-adiabatic transitions through a conical intersection between the ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ states. The bifurcation of the KER signals in our simulation is statistically reliable and our simulation directly predicts experimental KER images, which is beyond estimations using averaged AIMD trajectories26. Moreover, metastable $\text{CH}_{3}\text{I}^{2+}$ ionic states can be formed with small probability on one of the ionic surfaces near zero pump-probe delay.
This work reveals that dynamics on the ionic PES plays an important role on the delay-dependent KER signals, especially in the regime of short pump-probe delays which cannot be described by a simple 1D model. Our work also confirms that Coulomb explosion imaging technique can effectively be used for direct visualization of conical intersections in coordinate space, provided sufficient temporal resolution.
The method developed in this work is highly efficient in calculating trajectories with non-adiabatic effects, which can be applied to molecular dynamics simulations that highly rely on statistics. It can also be employed to identify reaction channels with small probabilities such as roaming. In the future, we plan to employ this method to investigate photo-isomerization processes such as bond rearrangement in diiodomethane and ring opening in thiophenone.

### v2

@@ 3.4 Coulomb explosion at short pump-probe delays
Fig.  6: Simulated kinetic energy release at different pump-probe delays ranging from 0 fs to 50 fs using about 20,000 PD-CE trajectories at $T=100$ K. The Coulomb explosion is simulated on the lowest ionic potential energy surface of $\text{CH}_{3}\text{I}^{2+}$ which corresponds to $\text{I}^{+}(^{3}\text{P}_{2})$ final dissociation limit. (a) Signals including all active states during photodissociation, which will lead to both the I and $\text{I}^{*}$ channels. Results using the 1D model in Eq. (8) are shown in solid yellow and black curves. Results assuming pure Coulomb interaction are shown in dashed yellow and black curves. (b) Signals only from the 7th adiabatic state during photodissociation, which correspond to the $\text{CH}_{3}+\text{I}$ channel. (c) Signals only from the 9th adiabatic state during photodissociation, which corresponds to the $\text{CH}_{3}+\text{I}^{*}$ channel at long distance. All signals are counted using a bin size $h_{t}=0.5$ fs and $h_{E}=0.01$ eV.
@@
At short pump-probe delays (t<50 fs), the $\text{CH}_{3}\text{I}$ molecule is in the intermediate states of photodissociation. Non-adiabatic transitions also take place at this time scale. To investigate the non-adiabatic dynamics in this regime, additional 20,000 PD-CE trajectories are calculated with initial conditions sampled at $T=100$ K and with a pump-probe delay step $\Delta t=0.1$ fs up to 50 fs. We have used a lower temperature for sampling the initial conditions than in the simulations shown above in order to have less overlap between the signal from the two dissociating states. The Coulomb explosion part is still simulated on the lowest PES corresponding to $\text{I}^{+}(^{3}\text{P}_{2})$ threshold. The resulting delay-dependent KER is shown in Fig. 6. A bifurcation feature is present between 10 fs and 20 fs in Fig. 6(a), which indicates the occurrence of non-adiabatic transitions. The photodissociation separates to two channels leading to I and I* products. The simulation also shows that the 1D model discussed above is a poor approximation at short pump-probe delays, because the C-I stretching is coupled with the $\text{CH}_{3}$($\text{CH}_{3}^{+}$) umbrella motion and the C-I bending at small C-I distances, which is not characterized by the 1D model. Not only does the 1D model predict a fake potential well on the $\text{I}^{+}(^{3}\text{P}_{2})$ ionic PES, but it also underestimates the KER. Fig. 6(b) shows that only a small fraction of trajectories (about 20 percent) undergo adiabatic passage and dissociate into the I channel. Most trajectories (about 80 percent) start to hop to the 9th adiabatic state near 12 fs and dissociate into the $\text{I}^{*}$ channel, as shown in Fig. 6(c).
@@
The bifurcation of KER signals has not been observed in CEI experiments so far. One reason is that the conical intersection is very close to the FC point and $\text{CH}_{3}$ is a very light fragment. Thus, non-adiabatic transitions occur within 20 fs after UV excitation, which requires pump and probe pulses with very high temporal resolutions to resolve. One option to alleviate this difficulty might be choosing a molecule with heavier fragments such as $\text{CF}_{3}\text{I}$. An alternative method to map the non-adiabatic transitions is using ATAS, which has been applied to methyl halide in previous works69, 62, 35.
@@
Fig.  7: Creation of metastable $\text{CH}_{3}\text{I}^{2+}$ dications. (a) Potential energy surface of $\text{CH}_{3}\text{I}^{2+}$ corresponding to the $\text{I}^{+}(^{1}\text{S}_{0})$ threshold, projected onto the $\alpha$ and $R$ plane. The energy difference between adjacent contours is 0.1 eV. 4,000 PD-CE trajectories are calculated at $T=100$ K with the CE dynamics occuring on this surface. The black triangles indicate the final geometries of 956 PD-CE trajectories that remain bound after propagating for over 2 ps on the ionic PES, and are thus counted as $\text{CH}_{3}\text{I}^{2+}$ signals. (b) Simulated $\text{CH}_{3}\text{I}^{2+}$ ion yield as a function of pump-probe delay.
@@
As mentioned before, only the ionic state associated with the $\text{I}^{+}(^{1}\text{S}_{0})$ threshold has a shallow well on the PES and can support a metastable $\text{CH}_{3}\text{I}^{2+}$ state, as shown in Fig. 7(a). But it should be noted that this channel has a rather small transition probability during the IR multi-photon ionization. In order to investigate the origin of $\text{CH}_{3}\text{I}^{2+}$ signals observed in the CEI experiment30, we calculate 4,000 PD-CE trajectories at very short pump-probe delays ($t<10$ fs), for which the Coulomb explosion is simulated on the $\text{I}^{+}(^{1}\text{S}_{0})$ ionic PES. 956 trajectories form $\text{CH}_{3}\text{I}^{2+}$ states and do not break up into ionic fragments after propagating for over 2 ps, The final geometries are shown on top of the PES in Fig. 7(a). The $\text{CH}_{3}\text{I}^{2+}$ signal has a maximum yield near 1 fs, and decays rapidly with the pump-probe delay, as shown in Fig. 7(b). We therefore conclude that the $\text{CH}_{3}\text{I}^{2+}$ signal detected in the CEI experiment30 is due to this shallow well on the ionic PES, and the observed enhancement of the $\text{CH}_{3}\text{I}^{2+}$ yield near zero pump-probe delay is also consistent with the pathway described here, but it is broadened in the experiment due to the cross correlation between the pump and the probe pulses. The internal states of $\text{CH}_{3}\text{I}^{2+}$ may be confirmed by photoelectron-photoelectron-photoion coincidence experiments.
@@ 4 Conclusion and outlook
In this work, we have employed an effective method to directly simulate the photodissociation followed by Coulomb explosion of methyl iodide with the inclusion of non-adiabatic effects. Potential energy surfaces of both the neutral molecule and the doubly charged cation are built upon high-level ab initio electronic structure calculations. More than 40,000 trajectories are calculated on both the neutral and ionic PESs to obtain a statistically reliable interpretation of the delay-dependent KER signals. The non-adiabatic effects are treated using a Landau-Zener surface hopping algorithm, which yields a branching ratio in overall agreement with previous theories and experiments. Our simulation of the delay-dependent KER at large pump-probe delays shows good agreement with a recent experiment using coincident ion momentum imaging techniques30, with the KER signals reflecting two reaction channels. At short pump-probe delays, the simulation shows a clear bifurcation of the KER signals near 12 fs, indicating the occurrence of non-adiabatic transitions through a conical intersection between the ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ states. The bifurcation of the KER signals in our simulation is statistically reliable, and our simulation directly predicts experimental KER observables with good accuracy, which is beyond the scope of previous simulations using averaged AIMD trajectories29. Moreover, the simulation shows that metastable $\text{CH}_{3}\text{I}^{2+}$ ionic states can be formed with small probability on one of the high-lying ionic surfaces near zero pump-probe delay.
This work reveals that dynamics on the ionic PES plays an important role when modeling experimental pump-probe observables such as the KER, especially in the regime of short pump-probe delays which cannot be described accurately by a simple 1D model. Our work also confirms that the Coulomb explosion imaging technique can effectively be used for direct visualization of conical intersections in coordinate space, provided sufficient temporal resolution.
The computational procedure used in this work is efficient in calculating trajectories with non-adiabatic effects as we build PESs beforehand using a streamlined interpolation scheme, which can be applied to molecular dynamics simulations that highly rely on statistics, e.g., to identify reaction channels with small probabilities such as roaming. In the future, we plan to employ this approach to investigate photo-isomerization processes such as bond rearrangement in diiodomethane and ring opening in thiophenone.